Problem 1-22 Cost Terminology: Contribution Format Income Statement [LO1-2, LO1-4, LO1-6 Miller Companys total sales are $240,000. The companys direct labor cost is $28,800, which represents 30% of its total conversion cost and 40% of its total prime cost, its total selling and administrative expense is $36,000 and its only variable selling and administrative expense is a sales commission of 5% of sales. The company maintains no beginning or ending inventories and its manufacturing overhead costs are entirely fixed costs. Required 1. What is the total manufacturing overhead cost? 2. What is the total direct materials cost? 3. What is the total manufacturing cost? 4. What is the total variable selling and administrative cost? 5. What is the total variable cost? 6. What is the total fixed cost? 7. What is the total contribution margin?
Answer #1

1.)

Total manufacturing overhead cost = 28,800/30%*70% = 67,200

2.)

Total direct material cost = 28,800/40%*60% = 43,200

3.)

Total manufacturing cost = 43200+28800+67200=139200

4.)

Total variable selling and administrative cost=240000*5%=12000

5.)

Total variable cost =43200+28800+12000= 84000

6.)

Total fixed cost = 67200+(36000-12000)=91200

7.)

Total contribution margin = 240000-84000 = 156000